# Robust brain age estimation from structural MRI with contrastive learning

### v1

@@ Abstract
Estimating brain age from structural MRI has emerged as a powerful tool for characterizing normative and pathological aging. In this work, we explore contrastive learning as a scalable and robust alternative to supervised approaches for brain age estimation. We introduce a novel contrastive loss function, $\mathcal{L}^{exp}$, and evaluate it across multiple public neuroimaging datasets comprising over 20,000 scans. Our experiments reveal four key findings. First, scaling pre-training on diverse, multi-site data consistently improves generalization performance, cutting external mean absolute error (MAE) nearly in half. Second, $\mathcal{L}^{exp}$ is robust to site-related confounds, maintaining low scanner-predictability as training size increases. Third, contrastive models reliably capture accelerated aging in patients with cognitive impairment and Alzheimer’s disease, as shown through brain age gap analysis, ROC curves, and longitudinal trends. Lastly, unlike supervised baselines, $\mathcal{L}^{exp}$ maintains a strong correlation between brain age accuracy and downstream diagnostic performance, supporting its potential as a foundation model for neuroimaging. These results position contrastive learning as a promising direction for building generalizable and clinically meaningful brain representations.
@@ 1 Introduction
Brain aging is a complex and individualized process, involving structural changes such as cortical thinning and white matter loss, which vary widely across the population. Accurately modeling these subject-specific trajectories is a long-standing challenge in neuroscience, with high clinical relevance for detecting and tracking age-related neurodegenerative diseases. Recent deep learning (DL) approaches have shown promise in building brain-age predictors that learn mappings from anatomical MRI to chronological age (Peng et al., 2021). However, these models typically require large-scale training data, which necessitates the aggregation of scans from multiple imaging sites and studies. A major challenge with multi-site data is that DL models, especially deep neural networks, tend to overfit to site-specific artifacts, such as differences in scanner hardware, acquisition protocols, or magnetic field strength. This phenomenon leads to degraded performance on unseen sites and severely limits real-world applicability (Glocker et al., 2019; Wachinger et al., 2021). As a result, the development of models that generalize well across heterogeneous sources remains an open and pressing problem. The OpenBHB challenge (Dufumier et al., 2022) was introduced as a benchmark to address this issue, offering a large-scale, multi-site dataset designed to evaluate brain-age prediction under real-world generalization constraints. Traditional models participating in this benchmark have largely relied on supervised regression objectives, such as mean absolute error (MAE) (Cole et al., 2017; Jónsson et al., 2019; Peng et al., 2021). However, these frameworks typically ignore the presence of site-specific confounds during training, which may affect the learned representations. To address this, we previously introduced a Contrastive Learning (CL) framework for brain-age regression (Barbano et al., 2023a), based on a novel similarity-driven loss function, $\mathcal{L}^{exp}$, which encourages biologically meaningful and site-invariant embeddings. In this extended work, we scale up our previous approach both in data and scope: we employ over 20,000 structural T1 scans from four large public datasets (OpenBHB, ADNI, ICBM, and OASIS-3), and conduct a broad empirical study to analyze generalization, site robustness, clinical validity, and downstream transferability. Our contributions are as follows:
@@
2 Results
Table 1: Brain age regression results. Single means that the models were trained and tested on each dataset separately; Merge means that the models were trained on the combined healthy samples. OpenBHB Challenge Results report balanced accuracy (BAcc), internal MAE, external MAE, and the composite score $\mathcal{L}_{c}$.
OpenBHB ADNI ICBM OASIS3 OpenBHB Challenge Results
Method Train MAE $R^{2}$ MAE $R^{2}$ MAE $R^{2}$ MAE $R^{2}$ BAcc Int. MAE Ext. MAE $\mathcal{L}_{c}\,(\downarrow)$
L1 Single 2.73 0.92 3.64 0.58 4.33 0.75 4.32 0.65 6.7 2.67 4.18 1.82
$\mathcal{L}^{y-aware}$ (Dufumier et al., 2021) 2.65 0.92 3.47 0.61 4.21 0.76 4.26 0.65 6.60 2.66 4.10 1.82
$\mathcal{L}^{threshold}$ (Barbano et al., 2023a) 2.66 0.93 3.36 0.63 5.27 0.57 4.18 0.66 5.73 2.95 4.10 1.74
$\mathcal{L}^{exp}$ (Barbano et al., 2023a) 2.66 0.93 3.44 0.62 4.11 0.77 4.22 0.66 5.1 2.55 3.76 1.54
L1 Merge 2.59 0.93 3.45 0.61 3.77 0.82 4.08 0.68 9.33 2.42 2.43 1.19
$\mathcal{L}^{y-aware}$ (Dufumier et al., 2021) 2.70 0.91 3.40 0.63 3.73 0.82 4.08 0.67 7.3 2.61 2.53 1.15
$\mathcal{L}^{threshold}$ (Barbano et al., 2023a) 3.11 0.85 3.43 0.62 3.83 0.80 4.11 0.66 6.0 2.90 2.21 0.95
$\mathcal{L}^{exp}$ (Barbano et al., 2023a) 2.77 0.91 3.39 0.63 3.84 0.81 4.05 0.68 5.40 2.71 2.25 0.93
(a) External MAE vs training size (scaling)
(b) Site bias vs training size. $\mathcal{L}^{exp}$ stays consistent.
Figure 1: Scaling contrastive brain-age models with data size. (a) External mean absolute error (MAE) decreases as training size increases; all methods benefit from more data, reaching a MAE between 2.21 and 2.53. (b) Balanced accuracy (B. Acc) in predicting scanner site increases for most methods as training size grows, indicating increased site bias. In contrast, $\mathcal{L}^{exp}$ remains stable and low across all sizes, demonstrating robustness to site-specific confounds.
(a) Age delta (BAG) distribution for HC, MCI, and AD classes across all datasets. The shaded areas represent the age delta distribution (histogram) for each class. The horizontal black line in the middle highlights the mean value (explicitly annotated for readability).
(b) ROC plots for using BAG as a proxy for detecting Alzheimer’s disease (AD) patients.
Figure 2: Using brain age gap (BAG) as a proxy for detecting neurodegeneration. (a) Age delta (BAG) distribution for healthy controls (HC), stable MCI (sMCI), progressive MCI (pMCI), and Alzheimer’s disease (AD) patients across all datasets. All methods consistently show increased BAG in pMCI and AD groups, indicating accelerated brain aging. (b) ROC curves evaluating the ability of BAG to separate HC and AD subjects in ADNI and OASIS datasets. Contrastive methods ($\mathcal{L}^{exp},\mathcal{L}^{threshold},\mathcal{L}^{yaware}$) match or slightly outperform the L1 baseline. (c) Longitudinal BAG trajectories (with $\mathcal{L}^{exp})$ on ADNI, showing that pMCI and AD subjects exhibit increased BAG over time, while HC and sMCI groups remain stable. This confirms that contrastive models capture disease progression dynamics.
2.1 Scaling training size improves generalization
In Tab. 1 we report results of brain age regression across the different datasets employed. We experiment with two pre-training schemes, namely, we first train and test on the same dataset (single), and then we train on the combined healthy population across all datasets and test on each test set separately (merge). As we can note from the results, the overall error decreases when pretraining on the larger dataset. We also benchmark our models on the public OpenBHB challenge, where we significantly improve the results in terms of generalization (external MAE). With $\mathcal{L}^{exp}$, we achieve state-of-the-art results in the challenge, scoring an impressive 0.93, compared to the previous SOTA of 1.54. By increasing the training set size, we are able to almost halve the external prediction error, meaning that the models generalize better. In Fig. 1(a), we show the scaling of models in terms of external MAE compared to the training set size. As we can see, all methods benefit from increasing the pre-training size. This is consistent with related literature (Schulz et al., 2024), showing that deep models follow exponential laws in brain age prediction from T1 structural MRIs.
2.2 $\mathcal{L}^{exp}$ is more invariant to site bias
By closely inspecting results in Tab. 1, we notice an interesting phenomenon: all models pre-trained on the merged dataset, other than $\mathcal{L}^{exp}$, exhibit a higher balanced accuracy (Bacc) in retrieving the acquisition site of the scan, meaning that as the training size gets larger, they become more biased towards the site. $\mathcal{L}^{exp}$ is the only method which preserve a final balanced accuracy below 6%. To illustrate these results in more detail, in Fig. 1(b), we show how the final balanced accuracy scales with respect to the size of the training data, in terms of acquisition sites and datasets. As we can observe from the results, our proposed loss $\mathcal{L}^{exp}$ shows a balanced accuracy almost constant at different training size, whereas most other methods are heavily influenced by the training composition, and tend to become more biased as diversity in the data increases. We hypothesize that this is due to the difference in the formulation of $\mathcal{L}^{exp}$ with respect to the other losses, such as y-aware. The presence of the exponential in the uniformity term of y-aware (Dufumier et al., 2021) may force some samples to be pushed apart from the anchor more than they should. This could result in a less “compact” latent space, that might capture more variance in the input data, such as noise from the site effect. On the other hand, $\mathcal{L}^{exp}$ may avoid this phenomenon as the samples are repelled according to the kernel value (Barbano et al., 2023a), achieving latent representations that are more invariant to such noise.
2.3 Contrastive Learning is a reliable estimator of accelerated aging
Next, we aim to assess whether CL is a reliable estimator to detect accelerated aging in subjects affected by cognitive impairment and neurodegenerative diseases such as Alzheimer’s Disease. Brain age is widely recognized as a biomarker of different brain disorders (Kaufmann et al., 2019), and many works proposed deep learning pipelines for brain age estimation (Cole et al., 2017; He et al., 2021). These models are typically trained with a supervised loss such as L1 and have shown robust performance in highlighting accelerated aging in affected patients. To assess whether our proposed CL framework also exhibits this property, we employ data from the Alzheimer’s Disease Neuroimaging Initiative (ADNI) with 1754 subjects and from OASIS-3 with 685 subjects for a total of 2439 subjects. The analysis is presented in Fig. 2. First, in Fig. 2(a), we show the distribution of brain age gap (BAG) - the difference between the model’s predicted age and the subject’s chronological age - across all groups healthy controls (HC), stable mild cognitive impairment (sMCI), progressive mild cognitive impairment (pMCI), and Alzheimer’s Disease (AD). The obtained results are mostly consistent among all methods, showing that contrastive losses are comparable to standard L1 supervised models. Furthermore, we notice that L1 exhibits a higher spread on the pMCI and AD classes, compared to contrastive methods. In Fig. 2(b) we report the ROC curves for detecting AD patients on ADNI and OASIS. Once more, the results show that contrastive learning methods are comparable to standard L1 supervised baselines, even with a slight edge (AUC of 0.82 vs 0.81 on ADNI and 0.73 vs 0.72 on OASIS). Lastly, we assess the longitudinal reliability of $\mathcal{L}^{exp}$ by calculating the change in BAG during follow-ups for ADNI subjects. As expected from Franke & Gaser (2012), we observe an increase of BAG for AD and pMCI patients, meaning an increased brain age during follow-ups; whereas HC and sMCI subjects remain stable.
2.4 Do better brain age models translate to better diagnostic models?
Finally, we also analyze whether more accurate brain age prediction models can translate to better diagnostic models. To do so, we finetune the models pre-trained on OpenBHB using ADNI data, and we plot the downstream accuracy in classifying AD and HC versus the original brain age MAE in Fig. 3. Consistently with related literature (Bashyam et al., 2020), we find that standard L1 supervised models do not benefit on average from lower age prediction error, showing a correlation between age prediction error and classification accuracy of $r=0.10$. In fact, Bashyam et al. (2020) note that tighter fitting models on age naturally produce brain-age estimates that offer little information beyond age, and moderately fitting brain age models obtain significantly higher differentiation in a transfer-learning setup. On the contrary, our results seem to suggest that models pre-trained using CL are able to learn complex and meaningful representations that can benefit downstream tasks. Specifically, $\mathcal{L}^{exp}$ shows an almost total correlation of $r=0.99$ between age prediction error and downstream accuracy, and exhibits the highest slope $m=0.34$ (meaning that the increase in accuracy is more rapid). This shows that CL is less prone to overfit the brain age prediction task. This is an important finding, as it may lay the foundation for robust pre-trained pipelines on neuroimaging data (e.g., foundation models).
Figure 3: AD classification accuracy (fine-tuning) vs brain age prediction error.
3 Discussion
Across all four datasets, training on the merged healthy cohort reduced the mean-absolute-error (MAE) relative to the single-dataset baseline (Tab. 1). The gains are most pronounced on the OpenBHB challenge, where the external MAE nearly halves on average. This mirrors the “scaling laws” recently reported for supervised models (Schulz et al., 2024), and shows that CL obeys similar dynamics. Importantly, Fig. 1(a) demonstrates that performance does not saturate in the range we explored, suggesting further room for improvement as larger public neuroimaging datasets become available.
$\mathcal{L}^{exp}$ mitigates site bias at scale
While every loss benefited from more data, only $\mathcal{L}^{exp}$ maintained a low balanced accuracy ($<6\%$) in predicting acquisition site after pre-training (Tab. 1, Fig. 1(b)). We hypothesize that the weighted repulsion term in $\mathcal{L}^{exp}$ prevents over-dispersion of negatives, yielding a latent space that captures biologically relevant variance while suppressing site-specific factors. By contrast, the fixed uniformity term in $y$-Aware and $\mathcal{L}^{threshold}$ can force hard negatives to drift apart excessively, leaving room for confounding factors to be learned. In this regard, L1 shows the strongest bias towards site. In practice, this robustness is crucial when models are deployed across centres that differ in hardware and protocol.
Contrastive pre-training is a reliable detector of accelerated ageing
All contrastive models replicated the well-established pattern of higher brain-age gaps (BAG) in pMCI and AD groups (Fig. 2(a)), also showing slightly lower variance than the L1 baseline. BAG-based ROC curves (Fig. 2(b)) equal or surpass the L1 reference (AUC 0.82 vs. 0.81 on ADNI; 0.73 vs. 0.72 on OASIS-3). Moreover, longitudinal trajectories on ADNI confirm that $\mathcal{L}^{exp}$ captures progressive neurodegeneration (Fig. 2(c)). Taken together, these results indicate that contrastive embeddings preserve clinically meaningful ageing signals rather than merely fitting chronological age.
For L1 models, tighter age fits do not predict better HC/AD separation (Pearson $r\!=\!0.10$), consistent with prior reports that over-fitting to age harms downstream transfer (Bashyam et al., 2020). Interestingly, the opposite seems to hold for $\mathcal{L}^{exp}$ ($r=0.99$; Fig. 3). We argue that contrastive pre-training encourages more robust representations that are relevant for disease, whereas strongly supervised objectives push all capacity toward age alone. This finding positions contrastive brain-age models as promising foundation models for a broad range of neuroimaging tasks.
Limitations and future work
First, although we aggregate over 20k scans, racial and ethnic diversity remains limited; future work should verify that the invariance of $\mathcal{L}^{exp}$ extends also to demographic shifts. Second, our analysis focuses on T1-weighted MRI; multi-modal extensions (e.g., diffusion or functional MRI) could further enrich the learned representations. Third, we treated the site-prediction task as a proxy for scanner bias; a more comprehensive assessment might include vendor, field strength, and pulse-sequence metadata.
Conclusions
Scaling contrastive pre-training with the proposed $\mathcal{L}^{exp}$ loss establishes new state-of-the-art performance on brain-age prediction and, crucially, yields embeddings that (i) are robust to scanner variation, (ii) capture accelerated ageing in neurodegenerative disease, and (iii) transfer monotonically to clinical classification tasks. These properties make CL a strong candidate for building large neuroimaging foundation models that can underpin future precision-medicine pipelines.
@@ 4.1 Experimental data
We use a collection of T1-weighted MRI scans comprising 7,045 different individuals, for a total of 20,647 images. The experimental data is gathered from different publicly available datasets: OpenBHB, ADNI, ICBM, and OASIS. The overall composition of the cohorts used in this study is presented in Tab. 2. OpenBHB (Dufumier et al., 2022) is a recently released dataset that aggregates healthy control (HC) samples from many public cohorts (ABIDE 1, ABIDE 2, CoRR, GSP, IXI, Localizer, MPI-Leipzig, NAR, NPC, RBP). Every scan comes from a different subject. We employ all the publicly available training data (N=3984). For testing, OpenBHB relies on two private test sets, one internal (with the same acquisition sites as the public training set), and one external (with different ones). For ADNI (Petersen et al., 2010) [^2] Data used in the preparation of this article were obtained from the Alzheimer’s Disease Neuroimaging Initiative (ADNI) database (adni.loni.usc.edu). , we include all phases ADNI-1, ADNI-2, ADNI-GO, and ADNI-3 in our study, amounting to 633 HC, 712 MCI (483 sMCI, 229 pMCI), and 409 AD patients. For ICBM (Mazziotta et al., 2001) [^3] Data used in the preparation of this work were obtained from the International Consortium for Brain Mapping (ICBM) database (www.loni.usc.edu/ICBM). and OASIS3 (LaMontagne et al., 2019), we include 622 and 685 subjects respectively, with the latter containing 88 AD cases.
@@
All the images underwent the same standard VBM preprocessing, using CAT12 , which includes non-linear registration to the MNI template and gray matter (GM) extraction. The final spatial resolution is 1.5mm isotropic and the images are of size 121 x 145 x 121. OpenBHB provides T1-weighted MRI scans acquired with 1.5T and 3T scanner. The data is already preprocessed with CAT12 and gray-matter volume is provided for download. For ADNI, we downloaded original T1-weighted images acquired with 1.5T and 3T scanners in the DICOM format, and we preprocessed them with the same pipeline of OpenBHB, using the brainprep package [^4] https://github.com/neurospin-deepinsight/brainprep, which is based on CAT12, in order to extract GM volumes. We included MRI scans acquired with the MP-RAGE protocol. For further details about the acquisition, we refer to the official documentation. [^5] https://adni.loni.usc.edu/methods/mri-tool/mri-analysis/. ICBM and OASIS3 provide T1-weighted MRI scans acquired with 1.5T and 3T scanners, and we preprocessed them with the same pipeline.

### v2

@@ Abstract
Estimating brain age from structural MRI has emerged as a powerful tool for characterizing normative and pathological aging. In this work, we explore contrastive learning as a scalable and robust alternative to L1-supervised approaches for brain age estimation. We introduce a novel contrastive loss function, $\mathcal{L}^{exp}$, and evaluate it across multiple public neuroimaging datasets comprising over 20,000 scans. Our experiments reveal four key findings. First, scaling pre-training on diverse, multi-site data consistently improves generalization performance, cutting external mean absolute error (MAE) nearly in half. Second, $\mathcal{L}^{exp}$ is robust to site-related confounds, maintaining low scanner-predictability as training size increases. Third, contrastive models reliably capture accelerated aging in patients with cognitive impairment and Alzheimer’s disease, as shown through brain age gap analysis, ROC curves, and longitudinal trends. Lastly, unlike L1-supervised baselines, $\mathcal{L}^{exp}$ maintains a strong correlation between brain age accuracy and downstream diagnostic performance, supporting its potential as a foundation model for neuroimaging. These results position contrastive learning as a promising direction for building generalizable and clinically meaningful brain representations.
@@ 1 Introduction
Brain aging is a complex and individualized process, involving structural changes such as cortical thinning and white matter loss, which vary widely across the population. Accurately modeling these subject-specific trajectories is a long-standing challenge in neuroscience, with high clinical relevance for detecting and tracking age-related neurodegenerative diseases. Recent Deep Learning (DL) approaches have shown promise in building brain-age predictors that learn mappings from anatomical MRI to chronological age (Peng et al., 2021). However, these models typically require large-scale training data, which necessitates the aggregation of scans from multiple imaging sites and studies. A major challenge with multi-site data is that DL models, especially deep neural networks, tend to overfit to site-specific artifacts, such as differences in scanner hardware, acquisition protocols, or magnetic field strength. This phenomenon leads to degraded performance on unseen sites and severely limits real-world applicability (Glocker et al., 2019; Wachinger et al., 2021). As a result, the development of models that generalize well across heterogeneous sources remains an open and pressing problem. The OpenBHB challenge (Dufumier et al., 2022) was introduced as a benchmark to address this issue, offering a large-scale, multi-site dataset designed to evaluate brain-age prediction under real-world generalization constraints. Traditional models participating in this benchmark have largely relied on standard supervised regression objectives such as L1 (Cole et al., 2017; Jónsson et al., 2019; Peng et al., 2021). However, these frameworks typically ignore the presence of site-specific confounds during training, which may affect the learned representations. To address this, we previously introduced a Contrastive Learning (CL) framework for brain-age regression (Barbano et al., 2023a), based on a novel similarity-driven loss function, $\mathcal{L}^{exp}$, which encourages biologically meaningful and site-invariant embeddings. In this extended work, we scale up our previous approach both in data and scope: we employ over 20,000 structural T1 scans from four large public datasets (OpenBHB, ADNI, ICBM, and OASIS-3), and conduct a broad empirical study to analyze generalization, site robustness, clinical validity, and downstream transferability. Our contributions are as follows:
@@
2 Related Works
Contrastive learning
When label information is not available, instance-based discrimination approaches (Chen et al., 2020) have gained considerable success, almost closing the gap with its supervised counterpart. Most recent research in CL has focused on designing the right data augmentations (Chen et al., 2020; Zhang & Ma, 2022), contrastive losses (Chen et al., 2020; Caron et al., 2021; Zbontar et al., 2021; Grill et al., 2020), memory banks (He et al., 2020; Dwibedi et al., 2021), prototypes (or clusters) (Caron et al., 2020; Li et al., 2020) typically for downstream classification, detection or segmentation tasks, leaving open the question for regression tasks. When weak labels are available (either discrete or continuous), kernel-based approaches have been explored for CL (Dufumier et al., 2021c, 2023; Sarfati et al., 2023; Dufumier et al., 2021b), using a re-weighting scheme to adapt the alignment term between samples with close weak labels. When continuous labels are available, ranking-based approaches (Zha et al., 2023) have recently gained traction by showing strong generalization results on regression tasks. Also recently, adaptive neighborhood selection in contrastive learning has been proposed (Träuble et al., 2024b, a) in the context of brain age regression using Magnetic Resonance Elastography (MRE) and demonstrated good generalization results for non-uniformly age distributed datasets.
Brain age estimation
Estimating the chronological age from brain imaging in healthy individuals has been a challenge for machine learning models for the past 10 years. The first traditional models (linear, SVM) were trained on handcrafted features extracted from structural MRI (cortical thickness, gray matter volume, surface area) (Cole et al., 2015; Franke & Gaser, 2012). Then, more complex deep learning models, such as CNNs (Cole et al., 2017) or Transformers (He et al., 2021), have been deployed to automatically extract brain features from raw imaging. These brain age models were used to derive a brain age gap correlated to common brain disorders such as schizophrenia, bipolar disorder, ASD, and ADHD (Kaufmann et al., 2019) and heritable (Cole et al., 2017; Jónsson et al., 2019). Nevertheless, recent studies (Dufumier et al., 2024) showed that DL models generalize poorly to new MRI sequences/acquisitions, and they still underperform on populations with different demographic characteristics. Additionally, it was shown that more recent Transformer-based architectures do not improve brain age models compared to classical CNNs, and they seem to extract similar concepts in their representations (Siegel et al., 2025). Finally, regarding the clinical utility of brain age models, a recent study (Tan et al., 2025) suggests that directly using the brain age gap as a biomarker for Alzheimer’s disease prediction might be suboptimal compared to models directly fitted for this task, in line with other studies (Schulz et al., 2025; Bashyam et al., 2020) suggesting that tightly-fitted brain age models do not necessarily transfer well on clinical tasks.
@@ 3.1 Experimental data
We use a collection of T1-weighted MRI scans comprising 7,045 different individuals, for a total of 20,647 images. The experimental data is gathered from different publicly available datasets: OpenBHB, ADNI, ICBM, and OASIS. The overall composition of the cohorts used in this study is presented in Tab. 1. OpenBHB (Dufumier et al., 2022) is a recently released dataset that aggregates healthy control (HC) samples from many public cohorts (ABIDE 1, ABIDE 2, CoRR, GSP, IXI, Localizer, MPI-Leipzig, NAR, NPC, RBP). Every scan comes from a different subject. We employ all the publicly available training data (N=3984). For testing, OpenBHB relies on two private test sets, one internal (with the same acquisition sites as the public training set), and one external (with different ones). It is worth noting that, thanks to the number of datasets gathered, OpenBHB provides one of the widest diversity in terms of population in neuroimaging data, including American, European, and Asian subjects. For ADNI (Petersen et al., 2010) [^2] Data used in the preparation of this article were obtained from the Alzheimer’s Disease Neuroimaging Initiative (ADNI) database (adni.loni.usc.edu). , we include all phases ADNI-1, ADNI-2, ADNI-GO, and ADNI-3 in our study, amounting to 633 HC, 712 MCI (483 sMCI, 229 pMCI), and 409 AD patients. For ICBM (Mazziotta et al., 2001) [^3] Data used in the preparation of this work were obtained from the International Consortium for Brain Mapping (ICBM) database (www.loni.usc.edu/ICBM). and OASIS3 (LaMontagne et al., 2019), we include 622 and 685 subjects respectively, with the latter containing 88 AD cases.
@@
All images underwent the same standard VBM preprocessing, using CAT12, which includes non-linear registration to the MNI template and gray matter (GM) extraction. The final spatial resolution is 1.5mm isotropic and the images are of size 121 x 145 x 121. OpenBHB provides T1-weighted MRI scans acquired with 1.5T and 3T scanner. The data is already preprocessed with CAT12 and gray-matter volume is provided for download. For ADNI, we downloaded original T1-weighted images acquired with 1.5T and 3T scanners in the DICOM format, and we preprocessed them with the same pipeline of OpenBHB, using the brainprep package [^4] https://github.com/neurospin-deepinsight/brainprep, which is based on CAT12, in order to extract GM volumes. We included MRI scans acquired with the MP-RAGE protocol. For further details about the acquisition, we refer to the official documentation. [^5] https://adni.loni.usc.edu/methods/mri-tool/mri-analysis/. ICBM and OASIS3 provide T1-weighted MRI scans acquired with 1.5T and 3T scanners, and we preprocessed them with the same pipeline.